OpenStudy (anonymous):

given the f(x)=-x^3+12x^2+60x-50 find 1st and 2nd derivatives

OpenStudy (welshfella):

find the derivative term by term use the algebraic formula if y = ax^(n) y' = an x^(n-1)

OpenStudy (welshfella):

so first derivative of -x^3 is -3 x^(3-1) = -3x^2

OpenStudy (welshfella):

- and for + 12x^2 it is +24 x can you continue/

OpenStudy (anonymous):

+24 x^2?

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

No the derivative of 12x^2 is 24x :- 12*2 x^(2-1) = 24x

OpenStudy (welshfella):

so far we have f'(x) = -3x^2 + 24x Now you have to find the derivative of 60x and -50

OpenStudy (anonymous):

how do u do it:/ sorry that I'm asking that

OpenStudy (welshfella):

using the formula